OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x − 9

OpenStudy (lgbasallote):

5x + 2= 4x - 9 subtract 4x from both sides 5x-4x -2 = -9 add 2 to both sides 5x - 4x = -9 + 2 combine..what do you get?

OpenStudy (anonymous):

7

OpenStudy (lgbasallote):

hmm close...which is bigger 9 or 2?

OpenStudy (anonymous):

9

OpenStudy (lgbasallote):

and what is the sign of 9? negative or positive?

OpenStudy (anonymous):

x is actually -7

OpenStudy (anonymous):

sorry -11

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

group like terms take the x over the equal sign then change the signe before it so positive 4x becomes - ...5x-4x=x the two u carried over becomes negatie -9-2=-11 ur welcome try some using this principle